If leah is 6 years older than sue, and john is 5 years older than leah, and the total of their ages is 41. Then how old is sue?
Answer by jorel1380(3719)

j=l+5
l=s+6
j=s+11
s+s+6+s+11=41
3s=24
s=8
Sue is 8. Leah is 14 and John is 19..